### v1

22email: antonino.nunnari@inaf.it 33institutetext: Department of Astronomy & McDonald Observatory, The University of Texas at Austin, 2515 Speedway, Austin, TX 78712, USA 44institutetext: ASI - Space Science Data Center, via del Politecnico snc, I-00133 Roma, Italy 55institutetext: Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden 66institutetext: Institute for Astronomy, University of Hawai’i at Manoa, Honolulu, HI 96822, USA 77institutetext: LMU München, Universitätssternwarte, Scheinerstr. 1, 81679 München, Germany 88institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 99institutetext: INAF – Osservatorio Astrofisico di Torino, via Osservatorio 20,10025 Pino Torinese (TO), Italy 1010institutetext: Department of Physics and Astronomy, Texas Christian University, TCU Box 298840 Fort Worth, TX 76129, USA 1111institutetext: INAF – Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, via Piero Gobetti 93/3, 40129 Bologna, Italy 1212institutetext: Dipartimento di Fisica, Sapienza Università di Roma, P.le A. Moro 5, I-00185 Roma, Italy 1313institutetext: INAF - Osservatorio Astronomico di Capodimonte, salita Moiariello 16, 80131, Naples, Italy 1414institutetext: Department of Science and Technology, Parthenope University of Naples, CDN-IC4, Naples, I-80143, Italy 1515institutetext: Dipartimento di Fisica e Astronomia, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: Department of Astronomy, School of Science, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan 1717institutetext: IAC- Instituto de Astrofísica de Canarias, Calle Vía Lactea s/n, E-38205 La Laguna, Tenerife, Spain 1818institutetext: Departmento de Astrofísica, Universidad de La Laguna, E-38206 La Laguna, Tenerife, Spain 1919institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, France
Classical Cepheids in the Galactic thin disk.
Classical Cepheids (CCs) have long been considered excellent tracers of the chemical evolution of the Milky Way’s young disk. We present a homogeneous, Non-Local Thermodynamical Equilibrium (NLTE) spectroscopic analysis of 401 Galactic CCs, based on 1,351 high-resolution optical spectra, spanning Galactocentric distances from 4.6 to 29.3 kpc. Using PySME with MARCS atmospheres and state-of-the-art grids of NLTE departure coefficients, we derive atmospheric parameters and abundances for key species tracing multiple nucleosynthetic channels (O, Na, Mg, Al, Si, S, Ca, Ti, Mn, Fe, Cu). Our sample–the largest CC NLTE dataset to date– achieves high internal precision and enables robust modeling of present-day thin-disk abundance patterns and radial gradients.
We estimate abundance gradients using three analytic prescriptions (linear, logarithmic, bilinear with a break) within a Bayesian, outlier-robust framework, and we also apply Gaussian Process Regression to capture non-parametric variations. We find that NLTE atmospheric parameters differ systematically from LTE determinations. Moreover, iron and most elemental abundance profiles are better described by non-linear behavior rather than by single-slope linear models: logarithmic fits generally outperform simple linear models, while bilinear fits yield inconsistent break radii across elements. Gaussian Process models reveal a consistent outer-disk flattening of [X/H] for nearly all studied elements. The [X/Fe] ratios are largely flat with Galactocentric radius, indicating coherent chemical scaling with iron across the thin disk, with modest positive offsets for Na and Al and mild declines for Mn and Cu.
Finally, Cepheid kinematics confirm thin-disk orbits for the great majority of the sample. Comparison with recent literature shows overall agreement but highlights NLTE-driven differences, especially in outer-disk abundances. These results provide tighter empirical constraints for chemo-dynamical models of the Milky Way and set the stage for future NLTE mapping with upcoming large spectroscopic surveys.
Key Words.:
Classical Cepheids – Galaxy formation and evolution – Spectroscopy – Chemical Abundances – …1 Introduction
Radial abundance gradients play a key role in constraining Galactic chemo-dynamical models (palla20; spitoni23; lemasle22). The radial trends observed in the present-day chemical abundances of the thin disk, when compared with theoretical predictions across different Galactocentric distances and stellar ages, allow us to constrain its chemical enrichment history (chiappini01). Furthermore, radial gradients and their steady flattening as a function of time soundly support stellar radial migrations in shaping the current metallicity distribution function of the thin disk (hou00; prantzos23). Moreover, they also provide an independent constraint on the star formation efficiency when moving from the innermost to the outermost disk regions (spitoni23).
Radial gradients across the thin disk have been investigated by using different stellar tracers: open clusters (friel02; magrini23; myers22; occaso24; otto25; dalponte25), planetary nebulae (Maciel13; stanghellini18), HII regions (fernandez17), and OB stars (daflon04; nieva12; braganca19). Classical Cepheids (CCs) have also been extensively used to investigate metallicity distributions and chemical gradients in the Galactic thin disk (yong06; szildi07; romaniello2008influence; pedicelli2010new; luck11; lucklambert11; lemasle2013galactic, lemasle2017detailed; genovali2013metallicity, genovali2014fine, genovali15; korotin14; martin15; andrievsky16; da2016neutron, da2022new, da2023oxygen; proxauf2018new; luck18; inno2019first; kovtyukh22; ripepi22; trentin23, Trentin24B, trentin24A, trentin25). Cepheids are radially pulsating variable stars that serve as excellent distance indicators, since their individual distances can be estimated with an accuracy of 1-3%. They also trace young stellar populations (bono2024cepheids). CCs are central helium-burning, intermediate-mass stars younger than a few hundred million years, with pulsation periods ranging from days to roughly one year. They are ubiquitous across the thin disk and the Galactic center (matsunaga11; bono2024cepheids).
Since CCs are typically distributed across the thin disk, their intrinsic colors are affected by uncertainties due to reddening corrections. This is the main reason why atmospheric parameters are estimated directly from the spectra. Several approaches have been proposed to address this issue. In particular, proxauf2018new and da2022new adopted the line depth ratio method (kovtyukh07) based on empirical calibrations of different element pairs to derive effective temperature (kovtyukh07; elgueta24). A new temperature scale for Galactic CCs based on a data-driven, machine-learning technique applied to observed spectra, was recently suggested by lemasle20 in which the flux ratios of different spectral features were tied to the effective temperatures derived using the infrared surface-brightness method (hanke18).
A significant fraction of abundance analysis of CCs is based on Local Thermodynamic Equilibrium (LTE) atmosphere models, such as those of castelli2004new and gustafsson2008grid. However, both theoretical predictions and empirical evidence show that LTE modeling can be prone to possible systematics (lind24; bergemann25). These systematics become more relevant when moving from dwarf stars to giants, and from metal-rich to metal-poor populations (e.g., thevenin1999stellar; idiart2000non; bergemann2012non; hansen2013lte; fabrizio2021use). Moreover, NLTE corrections are element- and line-dependent, with neutral lines more affected than ionized lines (kiselman2001nlte; collet2005effects; merle2011grid; tautvaivsiene2015gaia; duffau2017gaia).
Chemical abundances for CCs based on a NLTE approach have been provided by andrievsky16 for all their analyzed elements, by luck11 for CNO elements, by martin15 for Fe and O, and by korotin14 for O. However, in these investigations the stellar atmospheric parameters were derived by using an LTE approach.
Recently, amarsi2020galah provided NLTE departure coefficients for H, Li, C, N, O, Na, Mg, Al, Si, K, Ca, Mn and Ba. Additionally, NLTE departure coefficients are provided by mallinson24 for Ti, by amarsi22 for Fe, by caliskan25 for Cu and by amarsi25 for S. These departure coefficient are computed on a grid of 3756 1D MARCS model atmospheres (gustafsson2008grid) covering , , and dex. These comprehensive grids allow proper spectral modeling and NLTE analysis.
In this study, we apply a complete NLTE approach—deriving both atmospheric parameters and chemical abundances—for Galactic CCs by using high–resolution optical spectra. Although the statistical sampling in the inner (Galactocentric distance, ) and in the outer () disk is sparse, there is solid empirical evidence for radial gradients for most of the elements that have been investigated. More specifically, our dataset spans a wide range in Galactocentric distances (=5–29 kpc) and provides homogeneous NLTE abundances for light (O), odd-Z (Na, Al, Cu), (Mg, S, Si, Ca, Ti), and iron peak (Mn, Fe) elements. The current analysis is a stepping stone for exploiting thousands of new high-resolution spectra for CCs that will be collected from upcoming large spectroscopic surveys in optical (dejong16, 4MOST) and near-infrared (cirasuolo12, MOONS). This paper is organized as follows: in § 2, we present the spectral dataset and the line list we adopted for the spectral analysis. In § 3 we show the results concerning the atmospheric parameters and the chemical abundances, including the gradients, and the kinematic properties. The summary of the results, the discussion and a brief outline of the near future developments of this project are presented in § 4. The paper also includes several appendices. Appendix A presents the tables included in the paper, while Appendix B discusses the new distances of two CCs. Appendix C provides details of the adopted spectral synthesis, and Appendix D presents the validation of NLTE atmospheric parameters and chemical abundances.
2 Dataset and method
The current spectroscopic sample includes 401 CCs distributed across the thin disk, as shown in Fig. 1111NASA/JPL-Caltech/R. Hurt (SSC/Caltech), https://milkyway-plot.readthedocs.io/en/. Among them, 379 CCs have already been collected and discussed in da2023oxygen (hereinafter da2023oxygen). We added 66 High-Resolution (HR) spectra for 22 CCs (proprietary plus public archives): 10 with and 4 with .
2.1 Spectral sample
In the spectroscopic sample presented in da2023oxygen, individual heliocentric distances were estimated using the following methods: Gaia DR3 parallaxes (72%), W1-band Period-Luminosity (PL) relations (25%), K-band PL (2%), and J-band PL (1%). This ranking was adopted in da2023oxygen to favor purely geometrical distances over distances from PL relations. On the other hand, the ranking of PLs is W1-K-J to minimize the impact of extinction, which is less severe at redder passbands. The da2023oxygen sample includes 1,285 HR spectra and more than 80% have a signal-to-noise ratio per pixel greater than 100. For a comprehensive description of the da2023oxygen spectroscopic sample the reader is referred to da2022new and da2023oxygen.
The da2023oxygen dataset was complemented with 22 CCs for which 40 HARPS-N, 10 UVES and 16 ESPaDOnS spectra were available from public archives: 35 HARPS-N are proprietary spectra collected in the Stellar Population Astrophysics222The Stellar Population Astrophysics is an ongoing project based on a Large Program conducted at the Telescopio Nazionale Galileo (TNG) using the HARPS-N and GIANO B echelle spectrographs for about 74 nights from 2018 to 2021 (origlia19) program, while the other 5 were available in the TNG database and analyzed in ripepi21. The additional UVES and ESPaDOnS spectra have been analyzed in trentin23; Trentin24B, martin15, andrievsky16 and kovtyukh22. Details of the spectroscopic sample, such as the spectrographs, the resolution, and the spectral coverage, are provided in Table 1. We estimated the individual distances by using the same approach adopted by da2023oxygen: Gaia EDR3 parallaxes for 20 CCs and W1-band PL for 2 CCs. For two CCs (ASAS 181024–2049.6 and ASAS SN J065046.50–085808.7) we estimated a different distance from andrievsky16 and trentin23. The new distance estimates are discussed in more detail in the Appendix B.
2.2 Spectral synthesis method
We utilized a Python version of Spectroscopy Made Easy (SME, piskunov17) for spectral synthesis, i.e. PySME333pysme-astro=0.6.20, https://github.com/MingjieJian/SME (wehrhahn21). PySME generates synthetic spectra based on a given set of atmospheric parameters, specified spectral intervals, and spectral resolution. It determines the optimal values for the selected atmospheric parameters by fitting the synthetic spectra to the observed data, taking into account the data uncertainties. The free parameters in the fitting process can include one or more stellar parameters, specific elemental abundances, and parameters related to atomic transitions in the line list. PySME provides two uncertainties according to two different methods. The first is based on SME statistics, i.e. uses a metric based on the distribution of the derivatives for each free parameter, estimating the cumulative distribution function of the generalized normal distribution (piskunov17). The second is based solely on the least-squares fit. Further details on PySME are presented in Appendix C.
We adopted the line list presented in Table 2 to estimate the atmospheric parameters, including
Fe I, Fe II, Ti I and Ti II lines. We first tested this line list in a NARVAL
solar spectrum with resolution R68,000 and in HARPS and NARVAL spectra of
16 Gaia benchmark stars, both dwarfs and giants
(blancocuaresma14, casamiquela25). Spectral lines of CCs are broader with respect to dwarfs and to red giants because of extended convective envelopes and pulsation,
which are parameterized by higher microturbulence () and
macroturbulence () values (da2023oxygen; luck81, luck85; bersier96).
Consequently, spectral lines that show minimal blending in the Sun may experience more
significant blending effects in CCs. Fig. 22 shows the difference between the atmospheric parameters derived in this work and the reference values for the Sun and the benchmark stars provided by blancocuaresma14.
The atmospheric parameters are estimated according to the following steps:
Definition of the line masks, with information on the spectral coverage (segments), lines and continuum regions, and lines used for the RV adjustment;
Definition of the continuum and RV optimization, we optimized the RV and the continuum level in each segment (, , );
Definition of the atmosphere model, the abundance scale and the NLTE grids. We used the grid of MARCS model atmospheres (gustafsson2008grid), that PySME-package defines as ”marcs2012”. We used solar abundances from asplund09 and the most updated versions of the NLTE grids 444https://zenodo.org/records/3888394 (amarsi2020galah). In order to have supersolar [N/Fe] and subsolar [C/Fe], typical of stars that have experienced mixing episodes during their evolution such as CCs, we used A(N)=8.23 and A(C)=8.2 for [Fe/H]=0.0;
Spectral fitting in two different steps. In the first step, the free parameters are , , , and A(Ti), while in the second step the free parameters are , , and A(Ti).
We focus on the elements for which NLTE grids are available in PySME, that are C, N, O, Na, Mg, Al, Si, S, K, Ca, Ti, Mn, Fe, Cu, Ba (amarsi2020galah, amarsi22, amarsi25; mallinson24; caliskan25). The line list highlighted in Table 3 is built from Gaia-ESO database according to the two quality parameters and (further details in Appendix C). We used a NARVAL solar spectrum and three UVES spectra of Red Giant Branch (RGB) stars of M67, in order to validate our line list on giant stars for which the chemical abundances are well established (occaso24). In PySME, we use the abundances of the specific element and the macroturbulence as free parameters. The final [X/Fe] value is the median value of the [X/Fe] values of each line. Table 4 includes the [X/Fe] values of the calibrating stars.
We estimated the sensitivity of each spectral line to the respective chemical abundance, computed as the mean abundance variation over 10 CCs spectra, in response to changes in the atmospheric parameters. We varied of 50 K and 100 K, of 0.1 and 0.2 dex, of 0.05 and 0.1 dex, and of 0.2 and 0.4 km s-1. The results are listed in Table 5
3 Abundance Analysis
This section deals with the estimate of atmospheric parameters and the comparison with similar estimates available in the literature. Furthermore, we also discuss elemental abundances and their radial gradients with a comparison of the NLTE and LTE results. Finally, we present the kinematic properties of the sample.
3.1 Atmospheric parameters
We estimated the atmospheric parameters in LTE once the NLTE grids were turned off, leaving the rest of the code unchanged. Fig. 2 shows the difference of the atmospheric parameters and of iron and titanium abundances derived under NLTE and LTE assumption, with histograms. Effective temperature, surface gravity and microturbulence velocity are the most affected by the departures, while Fe and Ti display minimal variations, indeed the differences are, on average smaller than -0.03 dex. Although, the difference between NLTE and LTE abundances is modest and the standard deviation is, on average, smaller than 0.08 dex, the individual differences can be of the order of dex. Individual differences are in several cases larger than random errors, thus suggesting that the NLTE approach gives narrower metallicity distribution functions (smaller standard deviations).
Previous investigations addressing NLTE effects on the atmospheric parameters of FGK-type stars (see ruchti13; kovalev19) show that NLTE analyses generally yield higher , and values, particularly at low metallicities, as well as systematically higher microturbulent velocities of about for . Cepheids are supergiants and the is intrinsically much higher. The increase is supported by data plotted in Fig. 2, where is the only parameter that shows a statistically significant offset (), suggesting that microturbulence has a substantial impact on the determination of the other atmospheric parameters, which explains why we also find negative NLTE-LTE differences in , and . Further discussion and a comparison with previous studies is presented in Appendix D.
To constrain possible differences with similar investigations we also compared our NLTE atmospheric parameters and abundances with those provided by da2023oxygen by using the same spectra (Fig. 3).
The LTE atmospheric parameters provided by da2023oxygen were estimated by using pyMOOGi (adamow17), a Python version of the MOOG code (sneden73, sneden2002), atmosphere models from ATLAS (castelli2004new), and the ARES code (sousa07, sousa15) to measure the equivalent widths (EWs). In da2023oxygen, the authors used a large line list composed of 521 Fe I and Fe II lines based on different datasets and validated on calibrating Cepheids (proxauf2018new; da2022new). In contrast, our line list consists of 29 Fe I, Fe II, Ti I and Ti II, and was built in such a way as to be able to estimate atmospheric parameters via synthesis (see Sec 2.2 and Appendix C). Moreover, they derived by using the line depth ratio method (kovtyukh07; proxauf2018new), and , , and following the classical approach, i.e. consecutive iterations to reach ionization equilibrium for Fe I and Fe II lines and no trends in the Fe I abundances versus the EWs. We found positive trends in the difference of the atmospheric parameters, and indeed the slopes range from for the effective temperature (panel a) to for the surface gravity (panel b) and to for the microturbolence (panel c), but the mean differences are well within 1. The difference in iron abundance also shows a positive slope, which can be attributed both to differences in spectral analysis and to the fact that, as expected, the more metal-poor Cepheids appear even more metal-poor when analysed using a NLTE approach compared to LTE (bergemann2012non; hansen2013lte). However, the mean difference is quite small ().
Furthermore, we compared our iron abundance with those from previous studies, specifically we do have 101 Cepheids in common with Trentin24B and 209 Cepheids in common with luck18. The top panel of Fig. 4 shows that the comparison with Trentin24B covers a broad range in iron abundance ( dex) and the difference does not show any trend, while the comparison with luck18 (bottom panel of the same figure) covers a narrower range in iron abundance ( dex), but the difference displays a clear positive trend. The mean differences are quite small 0.02 0.18 dex (Trentin24B) and -0.080.11 dex (luck18), but the individual differences, once again, reach 0.4 dex.
The reader interested in a more detailed discussion concerning the atmospheric parameters is referred to appendix C.
3.2 Iron radial gradient
The Galactocentric distances covered by our sample range from 4.6 kpc to 29.3 kpc, with four CCs located in the outskirt of the thin disk ( kpc).
To investigate the iron radial gradient we adopted three different analytical fitting functions (Fig. 5) by taking into account errors on both axes: , , and a bilinear function. Moreover, we adopted a Bayesian inference framework to estimate model parameters and their uncertainties. Specifically, given data and parameters , the posterior probability is:
where is the prior and the likelihood. We work with the log-posterior . We sampled the posterior with the affine-invariant Markov Chain Monte Carlo (MCMC) sampler implemented in the emcee library to obtain parameter point estimates (MAP) and credible intervals (the 68% credible region from the marginal posterior). In particular, we adopted mixture models (foremanmackey14), for which the per-data-point likelihood is given by a weighted sum of a foreground (signal) and a background (outlier) component. Writing for the mixing fraction (the prior probability that a datum belongs to the background component), the contribution of datum is:
and the total log-likelihood is . The parameter therefore controls the relative weight of the background/outlier model and robustifies the inference against points that are inconsistent with the foreground model. In Gaussian implementations the components are typically normal distributions with their own means and variances (optionally including an extra “jitter” term added in quadrature to account for underestimated measurement errors). Numerical evaluation of should use stable log-sum-exp arithmetic to avoid underflow when one component is much smaller than the other. Convergence of the MCMC chains was verified via autocorrelation-time estimates and visual inspection of trace plots.
We provide the mean square error, the reduced chi-squared and the Akaike’s information criterion (AIC) scores for each fit. The AIC is a method for evaluating and comparing statistical models, which provides a measure of the quality of a statistical model’s estimate, taking into account both the goodness of the fit and the complexity of the model. It is defined as , where is the number of parameters of the model and is the maximized value of the total likelihood. Models with lower AIC should be preferred.
Data plotted in the top panel of Fig. 5 display quite clearly that the bilinear and the logarithmic fit reproduce quite well the radial gradient of CCs when moving from the innermost to the outermost disk regions. Note that the agreement for the bilinear fit is expected, since this fit has more degrees of freedom when compared with the linear and the logarithmic fit. However, the bilinear fit is also prone to possible systematics, since the edge of the two different intervals in Galactocentric distances is arbitrary. The maximized likelihood is obtained at a value of the ”knee” of (green dashed vertical line). In any case, the logarithmic fit is the most accurate analytical representation of the radial gradient, since the statistical parameters (MSE, reduced chi-squared, AIC score) attain their smallest values when compared with the other fits.
Fig. 6 shows the comparison between our linear iron radial gradient with those available in the literature, revealing a slope of -0.064 dex/kpc, which is steeper than the slopes provided by magrini23 by using open clusters and by luck18 by using CCs, and shallower than the slopes provided by Trentin24B based on CCs and by otto25 by using open clusters.
Fig. 7 displays the comparison of the current iron radial gradient with that of da2023oxygen in a plane. The slope of the current gradient is steeper (-1.53 vs. -0.91). The two radial profiles nearly overlap for kpc, but the current one becomes steeper in the outer disk. This difference is due to several reasons: the different spectral analysis method (including a different line list), the increased number of CCs at , which were not included in the da2023oxygen sample, and NLTE effects, since the outer disk Cepheids are the most metal-poor stars in the current sample.
To overcome possible problems in the choice of the adopted analytical function to describe the iron radial gradient, we decided to use the Gaussian Process Regression (GPR) to model the data without assuming a predefined functional form. The GPR is a Bayesian non-parametric method for modeling an unknown function in a regression setting, starting from a Gaussian process prior:
defined by a mean function and a covariance (kernel) function . Observational data (with noise ) induce a predictive posterior distribution for at new input points. This posterior yields both a predictive mean (the best estimate of the function) and a predictive variance (quantifying uncertainty in the estimate). The kernel function encodes assumptions about smoothness, length scales and correlation structure of the function. Hyper-parameters of the kernel are typically inferred by maximizing the marginal likelihood (or by Bayesian integration). Therefore, GPR adaptively fits complex, non-linear relationships while providing principled uncertainty quantification. In short, GPR offers a flexible and rigorous framework for regression when uncertainty estimates are required and when one prefers to avoid strong parametric assumptions on the functional form.
We employed GPR by using Gpy555https://gpy.readthedocs.io/en/deploy/. After testing different kernel functions, we adopted GPy.kern.Matern32 according to the distribution dispersion and different sampling across the x-axis. Furthermore, we optimized the variance and the length-scale of the kernel function.
Data plotted in Fig. 8 display several interesting features worth being discussed in more detail. The GPR fits quite well the distribution of CCs as a function of the Galactocentric distance and this outcome applies to both the inner and the outer disk Cepheids. The GPR shows a particular trend in the data across the Solar circle, a ”shoulder” at 8 kpc. Moreover, GPR also shows a steady change in the slope for . This evidence indicates that the linear fit fails to take account for the radial gradient variations of the CCs in the outer disk, thus suggesting that the change is an intrinsic feature when moving into the outer disk. The radial distribution of open clusters younger than 400 Myr collected by otto25 agrees quite well with the radial distribution of Cepheids. This agreement is expected, since CCs are stellar tracers younger than 300 Myr. Unfortunately, the range in Galactocentric distances covered by this sample is too small when compared with CCs, but the agreement is very promising.
3.3 Radial gradients of chemical abundances
Distributions of the chemical abundance ratios are shown in Fig. 9, together with overlaid Gaussian fits (dashed red lines). The present results indicate Solar abundance ratios, within 1, for four -elements (O, Mg, Si, S) and for one light metal (Al). In contrast, [Na/Fe] ratio shows a clear overabundance of +0.300.12 dex, while [Cu/Fe] appears underabundant, with a mean value of dex. The Na overabundance is expected, since Cepheids are intermediate-mass stars and the low–mass tail experiences the first dredge-up along the red giant branch (see Fig. C13 in bono2024cepheids). A similar Na overabundance was also reported by Trentin24B, who found a mean value of dex, whereas their [Cu/Fe] measurements yielded a mean value of . otto25, analyzing open clusters who derived abundances for both main-sequence and giant stars, obtained mean values of dex and dex. In passing, an overabundance in [Na/Fe] was also found by genovali15.
Fig. 10 shows the radial abundance gradient of ten out of the eleven investigated chemical elements as a function of , with blue solid lines representing the GPR models. Similar to iron, a clear flattening of the gradients at kpc is obtained for all the elements, but for oxygen. Unfortunately, O abundances are only limited to CCs located within 14 kpc, since our Oxygen-determination is based on the triplet at 7770 which is only covered by FEROS, STELLA and ESPaDOnS spectra in our spectroscopic dataset.
Flattening of the gradients is also supported by data plotted in Fig. 11, where the linear, logarithmic and bilinear fits are shown. The AIC scores support the logarithmic fit for Al, Si, S, Ca, Ti, Mn and Cu among the three models. Moreover, the linear fit is always the least suitable model of the three, while Na and Mg favor the bilinear model with respect to the logarithmic one, but with a different knee position (9.8 vs 15.3 kpc).
To further investigate the radial variation of the abundance gradients, Fig. 12 presents the different [X/Fe] ratios as a function of Galactocentric distance. The radial trends remain approximately constant for most of the investigated elements across the entire disk. The abundance ratio slopes are vanishing, suggesting that these elements follow the same radial behavior as iron. Three elements, however, deviate from this pattern: Al, Mn, and Cu. Mn and Cu display well-defined negative gradients (-0.0180.003, -0.0090.005 dex , respectively), while Al shows a positive gradient (+0.0120.003 dex ). The negative trend observed for Mn supports the behavior found for the iron-peak elements (Mn, Co, Ni) in OCs by otto25, whereas the nearly flat [X/Fe] ratios for O, Mg, Si, S, and Ca (all with slopes ¡0.01 dex ) further confirm the overall homogeneity of -element enrichment across the thin disk. genovali15 also reported nearly constant trends for Na, Al and Si. In contrast, they found mild evidence of positive slopes for Mg () and Ca (). However, their results were based on a smaller Cepheid sample, characterized by higher dispersion and a more limited spatial coverage.
Fig. 13 shows for the investigated elements the abundance ratios as a function of (P) along with a linear fit. CCs obey to a well defined Period-Age relation (bono05, and references therein). The individual age steadily increases when moving from long- to short-period Cepheids. The abundance ratio of all the elements remain approximately constant across the full age range of the Cepheids in the sample. Only O, Na, Si, Ti, and Cu show a weak positive trend, 0.110.20, 0.100.03, 0.070.02, 0.090.02, 0.080.05, respectively. For oxygen, the higher dispersion and the shorter age coverage may affect the result. In other words, age-related effects are weak or negligible in the radial gradients and abundance ratios presented above. genovali15 found a slight negative trend for Ca (-0.15), in contrast with the null slope obtained in our study. In da2023oxygen, they found a slope of 0.080.02 in the [S/Fe] vs plane, while we obtained a weaker positive trend of 0.040.02. This suggests that the overall consistency of the abundances in our sample, which reduces the dispersion, allows for a more accurate assessment of whether a trend is present.
In Fig. 14 the [X/Fe] vs [Fe/H] planes are shown. We emphasize the positive trend observed for Mn. Manganese is mainly produced in thermonuclear supernovae (Th-SNe), with yields that increase with metallicity (badenes08; kobayashi09), which explains the positive trend. Sodium, copper and aluminum are predominantly produced in core-collapse supernovae (CC-SNe); however, we observe a slight positive trend for Cu and a negative trend for Na and Al, reflecting their different metallicity dependences. kobayashi20 predict a constant or slightly negative trend for Na and Al, and a slight positive trend for Cu over the metallicity range we cover (), in agreement with our findings. Moreover, for Na and Al, the slight positive trend in [Na/Fe] and [Al/Fe] vs corresponds to a negative trend in the [Na/Fe] and [Al/Fe] vs [Fe/H] plane. Similarly, the negative trends of [Cu/Fe] and [Mn/Fe] with Galactocentric distance correspond to positive trends in the [Cu/Fe] and [Mn/Fe] vs [Fe/H] planes. For oxygen, magnesium, silicon, sulfur, calcium, and titanium, we expect a roughly constant trend in the metal-poor region of the sample and a decreasing trend at higher metallicities. This behavior arises because these elements are predominantly produced in CC-SNe, whereas metal-rich stars are increasingly affected by Th-SN enrichment, which reduces their [X/Fe] ratios. Since Si, S, Ca and Ti also receive contributions from Th-SNe, the negative trend in the metal-rich region is weak. In contrast, O and Mg are more affected and show a higher decreasing trend.
3.4 Radial gradients: LTE vs NLTE
We measured the chemical abundances under the LTE assumption and performed the logarithmic fit of the LTE abundance gradients. Fig. 15 reveals that for almost all the elements the NLTE slopes of the logarithmic fit are shallower. Only S and Cu show a similar slope, which suggests negligible NLTE effects for the lines employed for these two elements. This is expected for S, since we used an S I triplet that is weakly prone to NLTE effects (duffau2017gaia). NLTE have a minor impact on Cu I lines in stars with dex. shi18 found a mean abundance increase of about +0.07 dex when applying NLTE corrections, a result previously suggested by yan16.
The histograms in Fig. 16 illustrate the comparison of the NLTE and LTE abundances where, on average, lower NLTE abundances for O (-0.74 0.21 dex), Na (-0.110.08 dex), Al (-0.070.10 dex), Si (-0.080.15 dex) and Cu (-0.130.17 dex), and more centered distributions for Mg (-0.030.08 dex), S (-0.020.10 dex), Ca (0.060.09 dex) and Mn (0.020.14 dex). Our results indicate that the oxygen lines used in this study (OI 777 nm triplet) are strongly affected by NLTE, consistent with amarsi16, who reported corrections of about 0.6 dex or higher at solar metallicities. For the other elements, the NLTE corrections are more moderate and agree with the findings of amarsi2020galah. In appendix D we further discuss the NLTE abundances for Mn and O.
In Fig. 17, we compare the slopes of our chemical gradients with those reported in previous studies. A general agreement is observed in the inner disk for most elements, although a slight underabundance is found for sulfur and copper when comparing the GPR profiles to the linear trends from other works. Moreover, the GPR profiles provide additional information on the behavior of Na, Mg, Al, S, Ca, Ti, Mn and Cu, revealing a shoulder at the solar circle similar to that obtained in the iron GPR profile. The GPR profiles also suggest a change of slope around 14-16 kpc, resulting in a flatter gradient in the outer disk compared to the linear trends of previous studies, particularly for Na, Mg, Al, Si, S, Ca, Ti, Mn and Cu.
In Fig. 18, we emphasize the agreement with the da2023oxygen logarithmic profiles for oxygen and sulfur. According to da2023oxygen, the radial gradient of sulfur is steeper than that of iron, and sulfur is on average under-abundant compared to oxygen. This discrepancy points to the need for a revision of chemical evolution models, particularly regarding sulfur yields from massive stars, in order to reproduce observed trends in the outer Galaxy. However, our results confirm the general behavior of sulfur but show a profile that is more consistent with iron and the other elements, rather than reproducing the steeper trend reported by da2023oxygen.
3.5 Kinematic properties
The kinematic properties of CCs provide fundamental insights to constrain their true nature of thin disk stellar population. Their orbits were integrated by using the galpy code, a Python library dedicated to Galactic dynamics (bovy15). Positional parameters (RA, DEC, distance) and proper motions are mainly based on Gaia DR3. Individual distances have been discussed in more detail in § 2.1, while radial velocities are mainly based on current spectroscopic measurements. The galactic potential employed is MWPotential2014.
To improve the sampling across the four quadrants, the current sample was complemented with CCs analysed by luck18 and by Trentin24B for which homogeneous elemental abundances are available. We did not include the 16 CCs located at 3–5.6 kpc and analyzed by using high resolution NIR spectra by matsunaga23, because of possible systematics between NIR and optical spectroscopy. Moreover, the quoted samples have a sizable number of objects in common with the current sample and they have been adopted to move their chemical abundances into our metallicity scale.
Fig. 19 shows nine different kinematic planes with the orbital properties of CCs compared with a sample of Type II Cepheids (TIICs). These planes are based on kinematic diagnostics that are quite useful to characterize Galactic stellar populations, and in particular to identify stars that are associated to the the different Galactic components (lane22; bonifacio24; Bono et al. submitted). According to a Galactocentric cylindrical frame, , and are the radial, the tangential and the vertical velocity, while and are the angular momentum component along the Galactic Z-axis and the total orbital energy. , , and are the radial, the azimuthal, and the vertical actions, while is the total action, defined as . is the normalized angular momentum (circularity), defined as where is the angular momentum of a circular orbit with the same maximum binding energy. Finally, denotes the maximum height along the orbit above the Galactic plane. The bulk of the kinematic sample has cold orbits typical of thin disk stars, as stated by the coherent distribution in the 9 planes. In particular, in the Toomgre diagram (panel a) our sample is grouped in the bottom-right corner, characterized by and low perpendicular velocities, typical of the thin disk component. On the other hand, TIICs cover a broad range in transversal velocity and a fraction of them also attain negative transversal velocities suggesting that they are on retrograde orbits. The stark difference between the two samples is fully supported by the Lindblad plane (panel c) in which CCs are distributed along the tiny region typical of thin disk stars, while TIICs move from the region typical of hot stellar components with radial orbits and random motions. The same outcome applies to the action-diamond diagram (panel f), since CCs are grouped in the right corner, since the major contribution to the action is given by their . Furthermore, CCs show high circularities (, which means almost circular orbits and strong rotation), low eccentricities and low values (panels g, h, i). The TIICs display a more complex behavior, since they move from a warm stellar component more typical of thick disk stars to hot stellar components and to retrograde orbits. A more quantitative analysis will be provided in a forthcoming paper (Nunnari et al. in preparation).
The left panel of Fig. 20 shows the radial distribution of CCs projected onto the Galactic plane. The symbols are color-coded according to the azimuth angle . We note that stars with typically reach smaller Galactocentric radii than those at . This trend may arise from the intrinsic spatial distribution of stars in the sample (for example, due to the presence of spiral arms), from the selection function of our dataset, or from a combination of both effects. Data plotted in the central and right panel of the same figure display that for Galactocentric distances larger than 11-12 kpc the height above the Galactic plane steadily decreases. This finding is expected, since the variation is mainly caused by the Galactic warp. Specifically, it is known from previous observations that the Galactic midplane is bent downwards (i.e. Z¡0) in the portion of the disk covered by our dataset. The reader interested in a more detailed discussion concerning the Galactic warp is referred to chen19, skowron19, lemasle22 and poggio25.
It is worth mentioning that the change in the slope of the radial metallicity gradient occurs at Galactocentric radii where the Galactic warp begins to become significant (i.e. approximately 11-12 kpc), tentatively suggesting that the two aspects might be related. It is important to note, however, that the spiral arms and bar-driven radial mixing can also potentially generate variations in the radial metallicity gradient (see e.g., buder25).
The orbital properties (radial velocity and circularity) of these objects fully support their association to the thin disk. However, more photometric and spectroscopic data are required to address this issue on a more quantitative basis.
4 Summary and conclusions
The present study represents one of the most comprehensive NLTE spectral analyses performed on CCs spanning an unprecedented range in Galactocentric distance (4.6–29.3 kpc) and elemental coverage including light, alpha, odd-Z, and iron peak elements. Such an approach is crucial because CCs, as young (ages younger than a few hundred Myrs), luminous super-giant stars, sample the current chemical enrichment of the Galactic thin disk with minimal contamination from older populations. We used the largest high-resolution optical spectral sample, for which we estimated the atmospheric parameters and chemical abundances. We constrained the iron and the chemical gradients for O, Na, Mg, Al, Si, S, Ca, Ti, Mn and Cu.
Previous studies available in the literature have modeled the chemical gradients of the thin disk, by assuming either a linear (luck18; genovali2014fine; Trentin24B) or a logarithmic (da2023oxygen) function. The fits based on a linear function overestimate the observed gradient in the inner disk ( kpc) and underestimate the observed gradient in the outer disk ( kpc). The difference with previous investigations, based on a linear fit, is mainly due to the larger range in Galactocentric distances covered by the current sample. This means that we are dealing with 22 CCs with smaller than 6 kpc and 33 CCs with a larger than 15 kpc666The distance of the Cepheid ASAS J062939-1840.5 was estimated by using a mid-infrared (W1-band) PL relation and we found kpc. The distance of this object was also estimated by trentin23 using period-Wesenheit relations and they found kpc..
A flattening of the abundance gradient in the outer disk has been reported by several studies (e.g., luck18; kovtyukh22; da2023oxygen). A similar behavior has also been observed in open clusters (magrini23; otto25), where a bilinear fitting approach has been adopted. Nevertheless, the AIC test performed by otto25 indicated that introducing two additional parameters in a bilinear model does not significantly improve the fit compared to a simple linear relation. In any case, such modeling inherently implies a break at a certain Galactocentric radius, raising a theoretical question about the physical origin of this discontinuity. For our sample, the best-fitting bilinear model applied to the iron gradient (Fig. 5) suggests a break, or ”knee”, at kpc. This model provides a better fit than the linear one, as indicated by smaller residuals and a lower AIC score. However, when the same bilinear approach is applied to the other elements, the inferred knee positions vary considerably, ranging from 8 to 15 kpc. The bilinear model yields the lowest AIC score only for Mg compared to alternative models (Fig. 11), although in this case the knee occurs at kpc, a value inconsistent with the other elements and with previous studies.
In da2023oxygen, a large sample of 356 CCs, including four at kpc, was analyzed using a log–log relation for Fe, O, and S. In Fig. 5, we applied the same functional form to fit the iron gradient. As expected, this approach yields lower mean squared error, chi-squared, and AIC values compared to the linear and bilinear models, except in the case of Mg, as noted above. Furthermore, as shown in Fig. 7, the slope we obtain differs from that reported in da2023oxygen, likely due to differences in spectral analysis, such as the line list and the use of NLTE grids for estimating atmospheric parameters, and to the increased number of Cepheids located in the outskirt of the disk. Overall, the log-log fit appears to be the most reasonable among the three analytical models, although sodium and magnesium are better modeled by the bilinear fit (Fig. 11), suggesting that a single analytical function may not fully describe all gradients. For this reason, we also employed Gaussian Process Regression, which does not require any predetermined functional form. This modeling confirms that the radial abundance gradients in the thin disk generally flatten with increasing Galactocentric distance across all studied elements.
The flattening of the gradients is a feature that extends to extragalactic studies. Works on other disk galaxies have shown that, although the inner regions show a negative gradient, the radial distribution of metallicity flattens to a virtually constant value beyond the isophote radius. This behavior has been found in the disk galaxy M83 (bresolin09, bresolin16), and subsequently confirmed in NGC 1512 and NGC 3621 (bresolin12; kudritki14), where the oxygen abundance remains homogeneous and approaches an almost constant value in the outer regions. These observations suggest that flattening is a common feature of the metallicity gradient in spiral galaxy disks.
The current metallicity of CCs can be compared with that of B-stars located in the solar neighborhood. According to nieva12, early B-type stars within 500 pc show a mean iron abundance of 7.520.03 dex, consistent with the solar value. This result is based on a NLTE spectral analysis of 29 early B-type stars. In the present work, the mean iron abundance of the 11 CCs located within 500 pc of the Sun is 7.490.13 dex, in perfect agreement with nieva12. Differently, the 93 CCs located at the solar circle (i.e. all Cepheids with kpc), exhibit a slightly iron underabundance of 7.420.10 dex, but still consistent within 1.
Our results for chemical abundance ratios as a function of reveal that [X/Fe] ratios remain approximately constant across the entire thin disk. The near-flat distributions, with average values within 0.10 dex and minimal slopes, indicate that most elemental abundances scale closely with iron throughout the Galactic disk. This consistency suggests that these elements trace the gas distribution in the thin disk in a coherent manner. Slight overabundances in Al and Na, underabundance in Cu, and the small but measurable slopes for Mn, and Cu point to subtle variations, yet they do not compromise the overall trend of roughly constant abundance ratios. The negative slopes observed for Mn and Cu are in agreement with previous findings for iron-peak elements (otto25).
Chemical gradients from previous studies, such as genovali2014fine and da2023oxygen, indicate that the outer disk exhibits a higher dispersion. This effect can be partially attributed to the Galactic warp (kerr57; oort58; lemasle22; poggio25), a large-scale distortion of the disk that affects the estimated Galactocentric distances. In regions where the warp is pronounced, Cepheids appear closer in Galactocentric distance than they would in a non-warped disk. However, lemasle22 noted that the dispersion introduced by the warp is too small to fully account for the larger spread around the mean metallicity gradient reported by genovali2014fine and da2023oxygen. Instead, this increased dispersion likely reflects larger uncertainties associated with distant stars, whose spectra typically have lower signal-to-noise ratio. In our study, the gradients do not show a significant increase in dispersion (Fig. 10), in agreement with lemasle22. It is worth noting that our sample contains fewer Cepheids in the outer disk ( kpc), where the warp is stronger, than in the inner regions. Nevertheless, the trend visible in Fig. 20 appears to be a direct consequence of the Galactic warp. Similar trends have also been observed using larger kinematic samples of Cepheids (lemasle22, Fig. A.3; poggio25).
Results for the sample of the youngest open clusters (OCs) from otto25 overlaps with that of Cepheids. Sodium is an exception, since Cepheids are on average overabundant in this element. It was suggested by takeda13 and confirmed by dasilva15 in a comparison between giant and dwarf field stars, that this overabundance is mainly due to stellar evolution effects: the first dredge-up brings some sodium from the interior to the atmosphere, altering its initial chemical composition (ventura13; lagarde21). In contrast, the chemical abundances of OCs have been measured in both giant and dwarf stars, with the latter still unaffected by significant mixing processes. Comparisons with other stellar populations, such as open clusters, generally reveal similar radial trends for iron-peak and alpha elements, further supporting the homogeneity of chemical enrichment in the thin disk, at least in the region between 5 and 14 kpc. These findings are consistent with previous studies (magrini23; yong12; otto25).
This study opens multiple possibilities for future investigation. In addition to the natural need of increasing the sample size, especially in the outer regions of the disk, we emphasize the following points:
The chemical abundances of other elements, such as heavy elements produced by the s- and r-processes, have yet to be determined. As discussed in Trentin24B, some elements-such as Scandium, Zinc, and Zirconium-show a bifurcation at the low-metallicity tail. This anomalous behavior may arise from variations in spectral line analyses or from the presence of distinct stellar populations associated with different spiral arms. Since our sample includes Cepheids located in various spiral arms, a future study focusing on these elements could help clarify this issue. However, NLTE grids for these elements are not yet available in PySME.
In the inner disk, for kpc, a double sequence can appear, one that follows the logarithmic profile and the other showing a change in the slope. Note that a significant decrease in the metallicity of the young stellar population, confined to the central/inner regions, has been found in barred spiral galaxies such as NGC 1365 and M83 (sextl24, sextl25). Furthermore, andrievsky16 discussed a change in the slope at kpc, suggesting that it might be caused by a reduced star formation rate compared to the outer disk regions. To account for such a change of slope in the radial gradients, it has also been suggested that the inner disk regions have also been affected by dilution effects caused by the infall of more metal-poor gas (genovali15, and references therein). Further extensive sampling at small is crucial to investigate this feature of the gradient and its implications for the chemical enrichment of the Galactic bar.
The [X/H] gradients exhibit similar slopes among the studied elements, and the [X/Fe] ratios remain approximately constant across the entire range of explored. This behavior may indicate that these elements could serve as potential tracers of the underlying gas density profile. However, further validation is required to confirm these trends, and it remains to be established whether such a result can be extended to the disks of external galaxies.
The line list employed in this work serves as a tool for estimating atmospheric parameters and has been specifically adapted to optical spectra of Galactic Classical Cepheids. However, further investigation is required to extend the applicability of this method to the near-infrared (NIR) domain to fully exploit the NIR spectroscopic data and reach the more extinct regions. In addition, the metallicity range over which the selected lines have been tested should be broadened toward lower values, as the current analysis only covers the metallicity interval typical of the Galactic thin disk (-0.8 to +0.4 dex). This extension is essential to enable the application of the line list to more metal-poor Cepheids of Magellanic Clouds and the Local Group.
In conclusion, our results suggest that a single analytic function is not sufficient to model all the gradients. For this reason, we provided the behavior of each element gradient with a GPR modeling. However, when using an analytical function, a log-log profile should be preferred. Our results confirm that the radial abundance gradients in the thin disk flatten with radius across all studied elements. This is consistent with theoretical expectations of inside-out disk formation and continuous chemical enrichment, where inner regions exhibit higher metallicity and alpha-enhancements due to more rapid star formation. Nevertheless, the outer disk is still poorly sampled to give a definitive conclusion.
The adoption of NLTE modeling throughout, from atmospheric parameters to elemental abundances, represents an important step forward relative to prior studies that often combined LTE atmospheric parameters with NLTE corrections. The method employed in this study could be extended to data that come from large spectroscopic surveys (4MOST777https://www.4most.eu/cms/home/, dejong16) and from new instruments (MOONS@VLT888https://vltmoons.org/, cirasuolo12; MAVIS@VLT999https://mavis-ao.org/mavis/, mcdermid20), which will allow to increase sample sizes and extend spatial coverage, allowing for finer chemical gradient mapping, especially in regions that we are now poorly sampling such as the inner and outer disk, and improved constraints on Galactic formation models.
Acknowledgements.
Part of the research activities described in this report were carried out with contribution of the Next Generation EU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 - From Research to Business (M4C2), Investment Line 3.1 - Strengthening and creation of Research Infrastructures, Project IR0000034 – “STILES - Strengthening the Italian Leadership in ELT and SKA”, CUP C33C22000640006. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. We thank the support of the project PRIN MUR 2022 (code 2022ARWP9C) ‘Early Formation and Evolution of Bulge and HalO (EFEBHO)’ (PI: M. Marconi), funded by the European Union—Next Generation EU, and from the Large grant INAF 2023 MOVIE (PI: M. Marconi). This work has made use of the VALD database, operated at Uppsala University, the Institute of Astronomy RAS in Moscow, and the University of Vienna. H.J. acknowledges support from the Swedish Research Council, VR (grant 2024-04989). E.P. is supported in part by the Italian Space Agency (ASI) through contract 2018-24-HH.0 and its addendum 2018-24-HH.1-2022 to the National Institute for Astrophysics (INAF). J.M.O. acknowledges support for this research from the National Science Foundation Collaborative Grants (AST-2206541, AST-2206542, and AST-2206543).Appendix A Tables
Table 1 lists the properties of the spectroscopic sample. Lines used for atmospheric parameters, iron and titanium abundance estimation are listed in Table 2, while those used for chemical abundances are highlighted in Table 3. Table 4 shows the chemical abundances of the 3 RGB stars of M67 and the Sun. Finally, table 5 lists the values of for each spectral line computed as the mean abundance variation over 10 spectra, in response to changes in the atmospheric parameters.
| Spectrograph | N. of spectra | N. of stars | Resolution (R) | Coverage [] |
| HARPS1 | 231 | 29 | 115000 | [3790-6900] |
| SES2 | 400 | 64 | 55000 | [3860-8800] |
| FEROS3 | 339 | 161 | 48000 | [3500-9200] |
| UVES4 | 373 | 224 | 40000 | [3800-9400] |
| HARPS-N5 | 40 | 29 | 115000 | [3780-6910] |
| ESPaDOnS6 | 16 | 13 | 81000 | [3700-10500] |
1
High Accuracy Radial velocity Planet Searcher (HARPS, mayor03)
2
STELLA échelle Spectrograph (SES, strassmeier04; strassmeier10)
3
The Fiber-fed Extended Range Optical Spectrograph (FEROS, kaufer99)
4
Ultraviolet and Visual Echelle Spectrograph (UVES, dekker00)
5
High Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N, cosentino12)
6
Echelle SpectroPolarimetric Device for the Observation of Stars (ESPaDOnS, donati06)
| El | Ion. state | [] | gfflag | synflag | |
| Fe | 1 | 5217.389 | y | y | 3.211 |
| Fe | 1 | 5364.871 | y | y | 4.446 |
| Fe | 1 | 5569.618 | y | y | 3.417 |
| Fe | 1 | 5576.089 | y | y | 3.430 |
| Fe | 1 | 5638.262 | y | y | 4.220 |
| Fe | 1 | 5679.023 | y | y | 4.652 |
| Fe | 1 | 5705.464 | y | y | 4.301 |
| Fe | 1 | 5778.453 | y | y | 2.588 |
| Fe | 1 | 5855.076 | y | y | 4.608 |
| Fe | 1 | 5883.816 | y | y | 3.960 |
| Fe | 1 | 5916.247 | y | y | 2.453 |
| Fe | 1 | 5956.694 | y | y | 0.859 |
| Fe | 1 | 6056.005 | y | y | 4.733 |
| Fe | 1 | 6151.617 | y | y | 2.176 |
| Fe | 1 | 6165.360 | y | y | 4.143 |
| Fe | 1 | 6240.646 | y | y | 2.223 |
| Fe | 1 | 6246.318 | y | y | 3.603 |
| Fe | 1 | 6315.811 | y | y | 4.076 |
| Fe | 1 | 6336.823 | y | y | 3.686 |
| Fe | 1 | 6380.743 | y | y | 4.186 |
| Fe | 1 | 6494.980 | y | y | 2.404 |
| Fe | 2 | 4993.350 | u | y | 2.807 |
| Fe | 2 | 5264.802 | y | y | 3.231 |
| Fe | 2 | 6084.102 | u | y | 3.200 |
| Fe | 2 | 6149.246 | u | y | 3.889 |
| Fe | 2 | 6432.676 | y | y | 2.891 |
| Ti | 1 | 4913.613 | y | y | 1.873 |
| Ti | 1 | 4981.730 | y | y | 0.848 |
| Ti | 1 | 5219.702 | y | y | 0.021 |
| Ti | 1 | 5689.460 | y | y | 2.297 |
| Ti | 1 | 5978.541 | y | y | 1.873 |
| Ti | 1 | 6091.171 | y | y | 2.267 |
| Ti | 2 | 4874.009 | y | u | 3.095 |
| Ti | 2 | 5211.530 | y | u | 2.590 |
| Ti | 2 | 5381.022 | y | u | 1.566 |
| Ti | 2 | 5418.768 | y | y | 1.582 |
| El | Ion. state | [] | gfflag | synflag | |
| O | 1 | 7771.944 | 9.146 | ||
| O | 1 | 7774.166 | 9.146 | ||
| O | 1 | 7775.388 | 9.146 | ||
| Mg | 1 | 5711.088 | u | y | 4.346 |
| Mg | 1 | 6319.237 | u | y | 5.108 |
| Si | 1 | 5645.613 | y | u | 4.930 |
| Si | 1 | 5665.555 | y | u | 4.920 |
| Si | 1 | 5684.484 | y | u | 4.954 |
| Si | 1 | 5701.104 | y | y | 4.930 |
| Si | 2 | 6371.371 | y | u | 8.121 |
| Ca | 1 | 5260.387 | y | y | 2.521 |
| Ca | 1 | 5588.749 | y | y | 2.526 |
| Ca | 1 | 5590.114 | y | y | 2.521 |
| Ca | 1 | 5867.562 | y | y | 2.933 |
| Ca | 1 | 6166.439 | y | y | 2.521 |
| Ca | 1 | 6455.598 | y | y | 2.523 |
| Ca | 1 | 6471.662 | y | y | 2.526 |
| Ca | 1 | 6499.650 | y | y | 2.523 |
| S | 1 | 6743.540 | u | u | 7.866 |
| S | 1 | 6757.150 | u | u | 7.870 |
| Mn | 1 | 5394.683 | y | y | 0.000 |
| Mn | 1 | 5420.328 | y | y | 2.143 |
| Mn | 1 | 6021.771 | y | y | 3.075 |
| Al | 1 | 5557.063 | u | u | 3.143 |
| Al | 1 | 6696.023 | u | u | 3.143 |
| Al | 1 | 6698.673 | u | y | 3.143 |
| Na | 1 | 5688.205 | u | y | 2.104 |
| Na | 1 | 6154.226 | u | y | 2.102 |
| Na | 1 | 6160.747 | u | y | 2.104 |
| Cu | 1 | 5105.530 | y | u | 1.389 |
| Cu | 1 | 5218.195 | y | u | 3.817 |
| Cu | 1 | 5220.065 | y | u | 3.817 |
| Cu | 1 | 5700.170 | y | u | 1.642 |
| Target | Mg | eMg | Si | eSi | Ca | eCa | S | eS | ||
| 08512898+1150330 | 0.01 | 0.08 | -0.02 | 0.14 | -0.08 | 0.05 | 0.04 | 0.10 | ||
| 08511268+1152422 | -0.01 | 0.07 | -0.04 | 0.05 | -0.01 | 0.04 | 0.08 | 0.10 | ||
| 08512283+1148015 | 0.01 | 0.08 | -0.01 | 0.14 | -0.03 | 0.06 | -0.01 | 0.10 | ||
| Sun | 0.04 | 0.03 | 0.00 | 0.02 | 0.00 | 0.03 | -0.05 | 0.10 |
| Ti | eTi | Mn | eMn | Al | eAl | Na | eNa | Cu | eCu | |
| 08512898+1150330 | -0.07 | 0.05 | -0.14 | 0.02 | -0.09 | 0.07 | -0.08 | 0.06 | 0.02 | 0.08 |
| 08511268+1152422 | -0.04 | 0.05 | -0.10 | 0.05 | -0.07 | 0.08 | -0.11 | 0.05 | 0.00 | 0.01 |
| 08512283+1148015 | -0.09 | 0.06 | -0.14 | 0.03 | -0.09 | 0.06 | -0.08 | 0.03 | -0.07 | 0.06 |
| Sun | -0.04 | 0.04 | 0.00 | 0.01 | -0.01 | 0.02 | -0.03 | 0.05 | -0.04 | 0.02 |
| [] | [dex] | [dex] | [K] | [km s-1] | |||||||||||||
| X | -0.1 | -0.2 | +0.1 | +0.2 | -0.05 | -0.1 | +0.05 | +0.1 | -100 | -50 | +100 | +50 | -0.2 | -0.4 | +0.2 | +0.4 | |
| 4874.009 | Ti | 0.111 | 0.070 | 0.256 | 0.291 | 0.280 | 0.338 | 0.162 | 0.105 | 0.192 | 0.207 | 0.252 | 0.237 | 0.300 | 0.375 | 0.154 | 0.095 |
| 5105.530 | Cu | -0.002 | -0.007 | 0.009 | 0.017 | 0.048 | 0.092 | -0.034 | -0.074 | -0.074 | -0.037 | 0.095 | 0.046 | 0.033 | 0.063 | -0.017 | -0.041 |
| 5211.530 | Ti | 0.117 | 0.087 | 0.190 | 0.220 | 0.201 | 0.238 | 0.119 | 0.082 | 0.149 | 0.153 | 0.174 | 0.166 | 0.212 | 0.267 | 0.112 | 0.078 |
| 5218.195 | Cu | -0.012 | 0.001 | -0.006 | -0.010 | 0.055 | 0.125 | -0.072 | -0.126 | -0.088 | -0.043 | 0.062 | 0.027 | -0.004 | 0.034 | -0.018 | -0.014 |
| 5219.702 | Ti | 0.150 | 0.151 | 0.161 | 0.156 | 0.167 | 0.168 | 0.150 | 0.129 | 0.099 | 0.137 | 0.177 | 0.172 | 0.164 | 0.163 | 0.157 | 0.155 |
| 5220.065 | Cu | -0.057 | -0.162 | 0.173 | 0.249 | 0.227 | 0.321 | 0.009 | -0.165 | -0.009 | 0.036 | 0.316 | 0.183 | 0.121 | 0.124 | -0.013 | -0.075 |
| 5260.387 | Ca | -0.024 | -0.043 | 0.012 | 0.026 | 0.054 | 0.109 | -0.080 | -0.157 | -0.031 | -0.026 | 0.019 | 0.010 | -0.003 | -0.011 | -0.010 | -0.016 |
| 5381.022 | Ti | -0.223 | -0.239 | -0.138 | -0.129 | -0.135 | -0.125 | -0.178 | -0.200 | -0.165 | -0.159 | -0.143 | -0.148 | -0.129 | -0.113 | -0.201 | -0.239 |
| 5394.683 | Mn | -0.013 | -0.018 | 0.005 | 0.012 | 0.030 | 0.061 | -0.043 | -0.084 | -0.112 | -0.054 | 0.099 | 0.053 | -0.001 | -0.002 | 0.007 | 0.000 |
| 5418.768 | Ti | -0.039 | -0.072 | 0.043 | 0.078 | 0.052 | 0.093 | -0.033 | -0.073 | -0.016 | -0.005 | 0.030 | 0.020 | 0.088 | 0.137 | -0.061 | -0.119 |
| 5420.328 | Mn | -0.003 | -0.007 | 0.006 | 0.005 | 0.053 | 0.102 | -0.049 | -0.098 | -0.083 | -0.041 | 0.094 | 0.049 | 0.006 | 0.010 | -0.001 | -0.004 |
| 5557.063 | Al | -0.035 | -0.038 | -0.011 | -0.004 | 0.041 | 0.091 | -0.062 | -0.110 | -0.057 | -0.042 | 0.044 | 0.007 | -0.020 | -0.016 | -0.013 | -0.009 |
| 5588.749 | Ca | -0.005 | -0.006 | 0.005 | 0.014 | 0.060 | 0.120 | -0.066 | -0.127 | -0.066 | -0.036 | 0.067 | 0.032 | 0.142 | 0.304 | -0.130 | -0.227 |
| 5590.114 | Ca | 0.031 | 0.033 | 0.024 | 0.033 | 0.086 | 0.140 | -0.027 | -0.084 | -0.045 | -0.017 | 0.115 | 0.070 | 0.070 | 0.130 | -0.035 | -0.055 |
| 5645.613 | Si | 0.008 | 0.011 | 0.004 | 0.002 | 0.067 | 0.126 | -0.057 | -0.098 | -0.059 | -0.028 | 0.063 | 0.037 | 0.017 | 0.026 | -0.003 | -0.005 |
| 5665.555 | Si | 0.000 | 0.001 | -0.001 | 0.000 | 0.047 | 0.096 | -0.039 | -0.073 | -0.046 | -0.024 | 0.048 | 0.024 | 0.008 | 0.016 | -0.004 | -0.006 |
| 5684.484 | Si | -0.027 | 0.005 | 0.010 | 0.017 | 0.054 | 0.085 | -0.065 | -0.134 | -0.067 | -0.030 | 0.054 | 0.031 | 0.035 | 0.068 | -0.028 | -0.057 |
| 5688.205 | Na | -0.002 | 0.006 | -0.001 | 0.002 | 0.085 | 0.165 | -0.066 | -0.124 | -0.064 | -0.036 | 0.054 | 0.025 | 0.062 | 0.154 | -0.061 | -0.113 |
| 5700.170 | Cu | 0.005 | -0.031 | 0.008 | 0.028 | 0.127 | 0.250 | -0.153 | -0.306 | -0.103 | -0.065 | 0.123 | 0.059 | 0.019 | 0.023 | 0.013 | 0.005 |
| 5701.104 | Si | 0.007 | 0.006 | 0.002 | 0.001 | 0.051 | -0.012 | -0.032 | -0.072 | -0.034 | -0.014 | 0.046 | 0.025 | 0.011 | 0.019 | -0.002 | -0.006 |
| 5711.088 | Mg | 1.010 | 1.009 | 0.709 | 0.709 | 0.749 | 0.791 | 0.670 | 0.631 | 0.673 | 0.690 | 0.744 | 0.725 | 0.732 | 0.761 | 0.683 | 0.664 |
| 5867.562 | Ca | 0.018 | 0.021 | 0.007 | 0.003 | 0.057 | 0.093 | -0.018 | -0.042 | -0.020 | -0.008 | 0.061 | 0.035 | 0.010 | 0.013 | 0.016 | 0.024 |
| 6021.771 | Mn | 0.014 | 0.018 | 0.007 | 0.006 | 0.062 | 0.120 | -0.050 | -0.100 | -0.094 | -0.043 | 0.108 | 0.058 | 0.037 | 0.070 | -0.016 | -0.037 |
| 6154.226 | Na | 0.007 | 0.015 | -0.009 | -0.016 | 0.050 | 0.101 | -0.046 | -0.088 | -0.057 | -0.029 | 0.053 | 0.026 | 0.007 | 0.016 | -0.007 | -0.011 |
| 6160.747 | Na | 0.008 | 0.013 | -0.006 | -0.010 | 0.057 | 0.113 | -0.053 | -0.103 | -0.054 | -0.027 | 0.052 | 0.025 | 0.020 | 0.040 | -0.017 | -0.032 |
| 6166.439 | Ca | -0.002 | 0.031 | 0.008 | 0.016 | 0.048 | 0.107 | -0.033 | -0.094 | -0.035 | -0.031 | 0.067 | 0.033 | 0.031 | 0.064 | -0.026 | -0.022 |
| 6319.237 | Mg | 1.064 | 1.072 | 0.745 | 0.740 | 0.801 | 0.851 | 0.695 | 0.662 | 0.696 | 0.719 | 0.799 | 0.779 | 0.752 | 0.763 | 0.741 | 0.737 |
| 6371.371 | Si | -0.323 | -0.173 | -0.001 | 0.041 | -0.159 | 0.064 | -0.104 | -0.358 | 0.016 | -0.014 | -0.223 | -0.253 | -0.132 | 0.006 | -0.208 | -0.177 |
| 6455.598 | Ca | -0.006 | -0.013 | 0.016 | 0.027 | 0.073 | 0.137 | -0.054 | -0.133 | -0.054 | -0.024 | 0.072 | 0.037 | 0.016 | 0.029 | -0.013 | 0.003 |
| 6471.662 | Ca | 0.014 | 0.003 | 0.028 | 0.045 | 0.089 | 0.148 | -0.048 | -0.083 | -0.066 | -0.033 | 0.123 | 0.074 | 0.071 | 0.123 | -0.022 | -0.057 |
| 6499.650 | Ca | 0.032 | 0.046 | 0.017 | 0.022 | 0.071 | 0.124 | -0.031 | -0.081 | -0.037 | -0.016 | 0.089 | 0.059 | 0.046 | 0.084 | -0.000 | -0.018 |
| 6696.023 | Al | 0.001 | -0.001 | 0.000 | -0.005 | 0.073 | 0.138 | -0.066 | -0.126 | -0.076 | -0.039 | 0.082 | 0.041 | 0.009 | 0.017 | -0.005 | -0.011 |
| 6698.673 | Al | -0.003 | -0.004 | -0.001 | -0.007 | 0.054 | 0.114 | -0.063 | -0.110 | -0.073 | -0.037 | 0.066 | 0.029 | 0.004 | 0.008 | -0.009 | -0.013 |
| 6743.540 | S | -0.205 | -0.315 | 0.111 | 0.201 | -0.118 | 0.035 | 0.147 | -0.176 | -0.120 | 0.064 | -0.076 | 0.084 | 0.085 | -0.097 | -0.238 | -0.031 |
| 6757.150 | S | -0.200 | -0.173 | -0.101 | -0.081 | -0.094 | -0.049 | -0.063 | -0.300 | 0.016 | 0.008 | -0.166 | -0.133 | -0.083 | -0.014 | -0.086 | -0.094 |
Appendix B New distance estimates for two Cepheids
The targets and have been estimated respectively at a distance of 5.64 kpc (andrievsky16) and of 10.04 kpc (trentin23).
andrievsky16 applied the method of the colour excess determination based on the use of the calibrating relation from friedman11. Here EW(DIB) is the equivalent width of the 6613 diffuse interstellar band (DIB), which is also seen in Cepheid spectra. Using this relation, they found that the value for is equal to 1.41. Combining the approximate relation , the ”absolute magnitude–pulsational period” relation of gieren98, and the mean V magnitude, they found a heliocentric distance of 5641 pc, and kpc. Given the , consistent with the rest of the sample, we used the Gaia parallax to obtain a heliocentric distance of kpc and kpc ( table, bailerjones21). Given this value so different from that obtained by the other authors, we used the W1-band PL relation, calibrated from Galactic CCs by wang18, obtaining a distance of kpc, consistent with the value that is obtained by the parallax method.
trentin23 found a heliocentric distance of kpc for using PW relations calibrated on Galactic CCs by ripepi20 and on LMC CCs by ripepi22. This measurement is in contrast with that of skowron19, who found 7.60.3 kpc using mid-IR PL relations. In a coherent manner as described in Sec. 2.1, we found a heliocentric distance of kpc using the Gaia DR3 parallax, which is consistent with that of skowron19 within 2.
Appendix C Spectral synthesis details
The required steps for computing atmospheric parameters by using PySME are the following:
1.
Observed spectrum: with a specific spectral resolution.
2.
Atomic linelist: in a specified format, i.e. VALD.
3.
Atmosphere model: we utilized MARCS12 (gustafsson2008grid).
4.
NLTE grids: we used amarsi2020galah grids.
5.
Initial chemical abundances: we used those from asplund09.
6.
Initial set of atmospheric parameters.
7.
Spectral intervals (Masks) around the spectral lines used for parameter estimates with an identification of continuum and line regions.
PySME is designed to work in conjunction with the Vienna Atomic Line Database (VALD, ryabchikova15), which provides a comprehensive collection of atomic and molecular transition parameters of astronomical interest. In order to have good estimators we selected spectral lines with accurate atomic parameters. We updated the values for the lines used as estimators by referring to the Gaia-ESO database (heiter21). Gaia-ESO provides two quality parameters for each line: gfflag, which assesses the accuracy of , and synflag, which evaluates the blending resolution between closely spaced lines. These parameters are rated as ”y” (recommended use), ”n” (not recommended), or ”u” (undecided quality). For the line selection we need a relatively large number of estimators, possibly both neutral and ionized lines to be sensitive to the gravity variations. We selected Fe I, Fe II, Ti I, and Ti II lines with gfflag and synflag rated as ”y” for their high quality. This selection includes 21 Fe I lines, 2 Fe II lines, 6 Ti I lines, and 1 Ti II line. To expand the Fe II line list, we also included lines with gfflag rated as ”u”, adding 3 additional lines. Similarly, for the Ti II line list, we included lines with synflag rated as ”u”, adding 3 more lines. Fig. 21 displays an example of the fit of a Ti I line in a Cepheid spectrum. The complete line list is presented in Table 2.
We compared the atmospheric parameters derived from our method and the reference values for the Sun and 16 benchmark stars in Fig. 22.
We point out a limitation given by the atmosphere model used, since the atmospheric parameters are estimated at the edge of the grid for 22% of the spectra, as shown in Figure 23, where a ”staircase structure” artifact is in place. This implies that for the objects at the edge of the grid the surface gravity is in general overestimated, and/or the effective temperature is underestimated. As shown in Fig. 24, we compared our values of and with those of da2023oxygen to quantify this effect. We found an average difference of -0.20 dex in and 5 K in . While for the temperature the underestimation is completely negligible, for the surface gravity the average difference is comparable with the errors. For this reason, our final estimations of the surface gravity uncertainties take into account this effect.
A double check on possible trends among the different parameters has been performed, and in Figs. 25 the iron and titanium abundances are shown as a function of the atmospheric parameters.
Furthermore, to provide a more quantitative analysis of the differences between the NLTE and LTE approaches, we made use of the large number of HR spectra collected by our group for 20 calibrating Cepheids (da2022new). For these objects, the HR spectra cover the entire pulsation cycle, making them an excellent benchmark for validating spectroscopic approaches. The reason is twofold. a) Classical Cepheids with moderate to large amplitudes show variations along the pulsation cycle of approximately 1000 K in , about half a dex in , and a factor of two in (from 2 to 4 ). However, the chemical abundances for each object are the same. This makes them a robust laboratory for constraining the physical assumptions involved when transitioning from LTE to NLTE analyses. b) Classical Cepheids trace a loop in both the color-magnitude diagram and the spectroscopic colour-magnitude diagram (or Kiel diagram; vs. ; see kudritzki20, kudritzki24). As a consequence, during the pulsation cycle a Cepheid can reach, at a fixed surface gravity, two different effective temperatures: one on the rising branch and one on the declining branch. This provides a unique opportunity to test the impact of on NLTE analysis while keeping both surface gravity and chemical composition fixed.
We focused our analysis on three Cepheids for which we have very detailed phase coverage: , (fundamental mode; Fig. 26), (fundamental mode; Fig. 27), and (first overtone; Fig. 28). The bottom four panels of these figures show the differences in atmospheric parameters and iron abundances between the NLTE and LTE analyses. These panels reveal overall consistency between the two approaches, as no significant trends or systematic offsets are observed. The same conclusion holds when comparing the LTE results from da2022new with our current NLTE analysis (top four panels of the same figures).
Appendix D Validation of the NLTE parameters and abundances
We compared our NLTE atmospheric parameters and chemical abundances with previous studies that investigated the impact of NLTE analysis, focusing in particular on the behavior of NLTE-LTE differences as a function of metallicity. ruchti13 introduced an optimized methodology, termed NLTE-Opt, specifically designed to mitigate the systematic biases typical of standard 1D LTE analyses for FGK-type stars. In this approach, is determined from Balmer lines profiles, while and [Fe/H] are derived from the NLTE ionization balance of Fe. Fig. 29 shows the systematic offsets in stellar parameters as a function of metallicity based on standard LTE analysis when compared with NLTE-Opt estimates. These trends are generally attributed to the inability of the LTE Fe I excitation balance to provide an accurate scale. As shown by the black circles representing the results of ruchti13, these differences decrease moving toward higher metallicities (see also kovalev19). Our Cepheids (in blue) follow the same general behavior, displaying a good agreement with ruchti13, particularly in (panel a), (panel b) and (panel d). However, the panel c) of the same figure shows the difference in microturbulence differences, revealing an offset of 0.20–0.30 . Since Cepheids typically exhibit higher micro- and macroturbulent velocities, both the observed trends and the intrinsic scatter should be interpreted with caution. We also note that is the only parameter for which we find a statistically significant offset between the NLTE and LTE determinations (-0.150.11 , see Fig. 2). Finally, our fitting method simultaneously optimizes all atmospheric parameters, which means that the contribution of significantly affects the determination of the others.
We independently verified our NLTE abundances by applying the line-by-line NLTE abundance corrections provided by the MPIA database101010https://nlte.mpia.de/ (kovalev18). Specifically, we focused on Mn and O. The atomic model for Mn is the same as that used in our study from bergemann19, while for O, we used the atomic model from amarsi2018, while the MPIA database relies on bergemann21. We performed this comparison for five sample spectra whose atmospheric parameters are representative of the whole population. We downloaded the NLTE abundance correction from the MPIA database for Mn (5394.68 , 5420.33 and 6021.77 ) and O (7771.94 , 7774.17 and 7775.39 ). As shown in Fig. 30, the comparison indicates good agreement for all Cepheids, especially for Mn. For oxygen the agreement is also good (differences within 0.1 dex) for three Cepheids, while the remaining two show discrepancies of approximately 0.2 dex. This is well within our observational uncertainties. Finally, the decreasing trend observed in the plane vs has also been obtained in vasilyev19 for the 777.4 nm line with both 1D and 2D models, and previously observed in Cepheids in luck13.